OpenStudy (anonymous):

Jesse received two offers to join a music program. If he is enrolled in the program for 4 months, which program should he choose? First Program: 6 free songs for joining and 1 free song every month Second Program: 2 free songs for joining and 2 free songs per month for the first 3 months only (A) Not enough information (B) Either-He will get the same number of free songs from both (C) The second program (D) The first program

OpenStudy (anonymous):

D

OpenStudy (anonymous):

witch one

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

I need a explanation

OpenStudy (anonymous):

1st program- he gets 6 for joining and 1 every month for 4 months so 4x1=4 +6=10 2nd program- he gets 2 for joining and 2 every month for 3 months so 3x2=6 +2=8

OpenStudy (anonymous):

It's d yup

OpenStudy (anonymous):

6 + (1x4) =10